One year ago, sheela was four times as old as her daughter sakshi. Six years hence, sheela's age will exceed her daughter's age by 9 years. The ratio of the present ages of sheela and her daughter is

Answer by vicgonzerx(31)

let x=sheela (m0ther)----present age
y=daughter preent age
(x-1)=4(y-1)--------a year ago---eq 1
(x+6)=(y+6)+9-------6years hennce ----eq 2
x/y= required --------------?present
@ equation 1
x-1=4y-4
x=4y-3-----then substitute to eq 2
x+6=y+15
4y-3+6=y+15
3y=15-6+3
y=4-----------------daughter age
then solve sheelas age
x-1=4(4-1)
x=13
then...the ratio is x/y=13/4--------------------